If 10.0 g of a substance A decomposes for 100 min, the mass of A remaining unreacted is found to be 1.25g. What is the half life of this reaction if it follows first order kinectics?

The rate law for a reaction tha is first order in the disappearance of a single reaction, X.

Intergrated form of first order rate law
Answer #1

initial mass = 10.0 g

remaining mass = 1.25 g

time = 100 min

for first order :

rate constant = 1/t ln (Ao / At)

                       = 1/ 100 ln (10 / 1.25)

                  k = 0.0208 min-1

half - life = 0.693 / k

               = 0.693 / 0.0208

half - life = 33.33 min

Rate law = k [X]

rate = 1/ t ln [Ao / At]